# Holonomy of the Obata connection on 2-step hypercomplex nilmanifolds

### v1

@@ 1. $(1)$
    $\mathfrak{z}(\mathfrak{h})$ $=$ $\left\{(X,V):X\in\mathfrak{z}(\mathfrak{g})\cap\ker\rho,\;V\in\bigcap_{\begin{subarray}{c}Y\in\mathfrak{g}\end{subarray}}\ker\rho(Y)\right\},$
    $\mathfrak{h}^{1}$ $=$ $\left\{(X,V):X\in\mathfrak{g}^{1},\;V\in\rho(\mathfrak{g})\mathbb{R}^{4k}\right\}.$
@@ 2. $(2)$
    If $\mathfrak{g}$ is $l$-step nilpotent and
    $m_{\rho}=\min\{j:\rho(X_{1})\cdots\rho(X_{j})=0,\ \text{for all}\ \ X_{1},\dots,X_{j}\in\mathfrak{g}\},$
    then $\mathfrak{h}$ is $k$-step nilpotent, where $k=\max\{l,m_{\rho}\}$. In particular, if $\mathfrak{g}$ is $2$-step nilpotent and $\rho(X_{1})\cdot\rho(X_{2})=0$ for any $X_{1},X_{2}\in\mathfrak{g}$, then $\mathfrak{h}$ is $2$-step nilpotent.
@@ 3. $(3)$
  • Proof.
    The only non-trivial statement is the second, which follows from the identity
    $\operatorname{ad}_{(X,U)}^{k}=\left(\begin{array}[]{c|c}\operatorname{ad}_{X}^{k}&0\\ \hline\cr B_{U}^{k}&\rho(X)^{k}\end{array}\right),$
    where the operator $B_{U}^{k}:\mathfrak{g}\to\mathbb{R}^{4k}$ is given by
    $B_{U}^{k}Y=\left(\sum_{l=1}^{k}(-1)^{l}\binom{k}{l}\rho(X)^{k-l}\rho(Y)\rho(X)^{l-1}\right)U.$
New examples of Obata non flat 2-step nilpotent hypercomplex Lie algebras can be obtained by applying the semidirect product construction, as we show next.
@@
We note that the simply connected Lie groups corresponding to the Lie algebras from examples 4.6 and 4.7 admit lattices, according to [Mal].
@@ 5 $\mathbb{H}$-solvability of $2$-step nilpotent Lie algebras
We recall the following definition.
Definition 5.1 ([Gor1]).
Let $\mathfrak{g}$ be a nilpotent Lie algebra endowed with a hypercomplex structure $(J_{1},J_{2},J_{3})$. Define inductively $\mathbb{H}$-invariant Lie subalgebras:
$\mathfrak{g}_{k}^{\mathbb{H}}:=\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{1}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{2}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]+J_{3}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],$
where $\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g}$. Note that $\mathfrak{g}_{1}^{\mathbb{H}}:=\mathbb{H}\mathfrak{g}^{1}=\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}+J_{2}\mathfrak{g}^{1}+J_{3}\mathfrak{g}^{1}$.
A hypercomplex nilpotent Lie algebra $\mathfrak{g}$ is called $\mathbb{H}$-solvable if the following sequence eventually becomes zero for some $k\in\mathbb{N}$:
@@
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a nilpotent hypercomplex Lie algebra. Then for any $k\in\mathbb{N}$, $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$ and it is an ideal in $\mathfrak{g}_{k-1}^{\mathbb{H}}$.
  • Proof.
    Since $\mathfrak{g}^{1}\subset\mathfrak{g}^{1}_{\mathbb{H}}$, $\mathfrak{g}^{\mathbb{H}}_{1}$ is clearly a subalgebra. By induction, assume that $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra. Then
    $\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k}^{\mathbb{H}}]=[\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}],\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]].$
    Since $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra $\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}_{k-1}^{\mathbb{H}}$. In particular
    $\mathfrak{g}_{k}^{\mathbb{H}},\mathfrak{g}_{k}^{\mathbb{H}}]\subset[\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}}]=[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\subset\mathfrak{g}^{\mathbb{H}}_{k}.$
    Since $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra for each $k$, we have:
@@
Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra endowed with a complex structure $J$. It was shown by Salamon [Sal, Corollary 1.4] that the ideal $\mathfrak{g}^{1}+J\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$. This naturally raises the question of whether an analogous property holds for nilpotent Lie algebras endowed with a hypercomplex structure.
@@
The conjecture holds true for Lie algebras of hypercomplex nilmanifolds with flat Obata connection [Gor2] and for nilpotent Lie algebras with abelian hypercomplex structure [AV, Proposition 4.5]. Here we prove the conjecture for $2$-step nilpotent Lie algebras.
@@
  • Proof.
    To prove Theorem 5.4, we show that the $\mathbb{H}$-solvable series (5.1) terminates at 0. In particular, this implies that for any $k\in\mathbb{N}_{\geq 1}$, $\mathfrak{g}_{k}^{\mathbb{H}}\subsetneq\mathfrak{g}$, proving that $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of the Lie algebra $\mathfrak{g}$.
@@
    By Lemma 2.4,
@@
    Since by Equation 2.5 $J_{\alpha}[J_{\alpha}\mathfrak{g}^{1},J_{\beta}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}$, we may write:
@@
    We claim that $J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]$. Using the quaternionic identities and Equation 2.5 with $J_{1}$, we have:
@@ 6 Holonomy of $2$-step hypercomplex nilmanifolds
We use Theorem 5.4 to prove a structural result about the holonomy algebra of the Obata connection on $2$-step nilpotent Lie groups endowed with invariant hypercomplex structures.
@@
In particular, it is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$ with trivial product, that is, for any $A,B\in\mathfrak{hol}(\nabla),\ A\cdot B=0$.
  • Proof.
@@
        $J_{2}[\cdot,J_{2}[\cdot,\cdot]]=[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, where the equality follows by Equation (2.5),
@@
        $J_{3}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows by Equation (2.5),
@@
        $J_{1}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows again by Equation (2.5).
      We set $\mathfrak{h}:=\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. It is not difficult to note that $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant and $\mathfrak{h}\subset\mathfrak{g}_{1}^{\mathbb{H}}$.
@@
        $Y,X]+J_{1}[J_{1}Y,X]=[Y,X_{2}]+J_{1}[J_{1}Y,X_{2}].$
        Using that $X_{2}\in J_{1}\mathfrak{g}^{1}$, $X_{2},Y]=-J_{1}[X_{2},J_{1}Y]$ by Equation (2.5). Therefore,
        $Y,X]+J_{1}[J_{1}Y,X]=-J_{1}[J_{1}Y,X_{2}]+J_{1}[J_{1}Y,X_{2}]=0.$
@@
      where we used twice Lemma 6.2 and the $\{J_{\alpha}\}$-invariance of $\mathfrak{h}$.
    • Step 3: Expression for the curvature endomorphisms and $\nabla$.
      By Theorem 5.4, the ideal $\mathfrak{g}_{1}^{\mathbb{H}}$ is a proper ideal of $\mathfrak{g}$ containing $\mathfrak{h}$. Therefore, we may write $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{g}_{1}^{\mathbb{H}}$, where $\mathfrak{k}_{1}\neq 0$ is any $\{J_{\alpha}\}$-invariant complement of $\mathfrak{g}_{1}^{\mathbb{H}}$. Similarly, we decompose $\mathfrak{g}_{1}^{\mathbb{H}}=\mathfrak{h}_{1}\oplus\mathfrak{h}$, where $\mathfrak{h}_{1}$ is any $\{J_{\alpha}\}$-invariant complement of $\mathfrak{h}$ inside $\mathfrak{g}_{1}^{\mathbb{H}}$ (such complements always exist, as we can consider the orthogonal complements with respect to a fixed hyperHermitian metric). We note that $\mathfrak{h}_{1}$ could be trivial; nonetheless, the proof proceeds in exactly the same manner without any differences.
      With respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$, using Step $1$ and Step $2$, we get that for any $X,Y\in\mathfrak{g}$,
      $R(X,Y)=\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}$.
      We have the following two Lemmas:
      Lemma 6.3.
      Let $X,Y\in\mathfrak{g}$. Then $\nabla_{X}Y\in\mathfrak{g}_{1}^{\mathbb{H}}$.
      • Proof.
        Straightforward by Formula (2.2). ∎
      Lemma 6.4.
      Let $X\in\mathfrak{g}$. Then $\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}}\subset\mathfrak{h}$.
      • Proof.
        Since $\nabla_{X}J_{\alpha}=J_{\alpha}\nabla_{X}$ and $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant, it suffices to prove that $\nabla_{X}\mathfrak{g}^{1}\subset\mathfrak{h}$. Applying Formula (2.2), we see that for any $Y\in\mathfrak{g}^{1}$:
        $\nabla_{X}Y=\frac{1}{2}\big{(}-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\big{)}.$
        By Step 1, the first term lies in $\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, while the second term lies in $J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. The proof follows by the definition of $\mathfrak{h}$. ∎
      By the previous Lemmas we get that with respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$:
      $\nabla_{X}=\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr\ast&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}$. (6.2)
@@
      We recall the following version of the Ambrose-Singer Theorem for left invariant connections.
@@
      [Al] Let $\nabla$ be a left invariant linear connection on the Lie group $G$, and let $\mathfrak{g}$ denote the Lie algebra of $G$. Then the holonomy algebra $\mathfrak{hol}(\nabla)$, based at the identity element $e\in G$, is the smallest subalgebra of $\mathfrak{gl}(\mathfrak{g})$ containing the curvature endomorphisms $R(X,Y)$ for any $X,Y\in\mathfrak{g}$, and closed under commutators with the left multiplication operators $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$.
      Let us consider $\text{span}\{R(X,Y)\ |\ X,Y\in\mathfrak{g}\}$. As already observed in Step $3$, with respect to the decomposition $\mathfrak{g}=\mathfrak{k}_{1}\oplus\mathfrak{h}_{1}\oplus\mathfrak{h}$,
      $\text{span}\{R(X,Y)\ |\ X,Y\in\mathfrak{g}\}\subset\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}\in\mathfrak{sl}(n,\mathbb{H})\right\}.$
      We set $\mathfrak{a}:=\left\{\begin{pmatrix}\begin{array}[]{c|c|c}0&0&0\\ \hline\cr 0&0&0\\ \hline\cr\ast&\ast&0\\ \end{array}\end{pmatrix}\in\mathfrak{sl}(n,\mathbb{H})\right\}$. We point out that $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
      Let $A\in\mathfrak{a}$ and let $X\in\mathfrak{g}$. Using the expression of $\nabla_{X}$ given in Step 3, the commutator
@@
      Since $\mathfrak{a}$ is closed under commutators with $\nabla_{\mathfrak{g}}$, by Theorem 6.5 we get that $\mathfrak{hol}(\nabla)\subset\mathfrak{a}$. In particular, since $\mathfrak{a}$ is abelian with trivial product, Theorem 6.1 is proved.
@@
Let $(\mathfrak{g},\{J_{\alpha}\})$ be the $2$-step nilmanifold hypercomplex Lie algebra constructed in Example 4.3 and let us consider the corresponding $2$-step nilpotent hypercomplex Lie group $G$. As already observed, the Obata connection $\nabla$ on $G$ is not flat. A direct computation shows that the holonomy algebra $\mathfrak{hol}(\nabla)$ is given by

### v2

@@ 1. $(1)$
    $\mathfrak{z}(\mathfrak{h})$ $=$ $\left\{(X,V)\mid X\in\mathfrak{z}(\mathfrak{g})\cap\ker\rho,\;V\in\bigcap_{\begin{subarray}{c}Y\in\mathfrak{g}\end{subarray}}\ker\rho(Y)\right\},$
    $\mathfrak{h}^{1}$ $=$ $\left\{(X,V)\mid X\in\mathfrak{g}^{1},\;V\in\rho(\mathfrak{g})\mathbb{R}^{4k}\right\},$
    where $\rho(\mathfrak{g})\mathbb{R}^{4k}=\mathrm{span}\{\rho(X)V\mid X\in\mathfrak{g}\ \text{and}\ V\in\mathbb{R}^{4k}\}$.
@@ 2. $(2)$
    If $\mathfrak{g}$ is $l$-step nilpotent and the set
    $\{j\in\mathbb{N}\mid\rho(X_{1})\cdots\rho(X_{j})=0,\ \text{for all}\ \ X_{1},\dots,X_{j}\in\mathfrak{g}\}$ (4.4)
    is non-empty, then $\mathfrak{h}$ is $r$-step nilpotent, with $r=\max\{l,m_{\rho}\}$, where $m_{\rho}$ is the minimum of the set defined in (4.4). In particular, if $\mathfrak{g}$ is $2$-step nilpotent and $\rho(X_{1})\cdot\rho(X_{2})=0$ for any $X_{1},X_{2}\in\mathfrak{g}$, then $\mathfrak{h}$ is $2$-step nilpotent.
@@ 3. $(3)$
The only nontrivial points are the second equality in (1) and statement (2).
The first claim follows from the computation
$\mathfrak{h}^{1}=[\mathfrak{h},\mathfrak{h}]=[(\mathfrak{g}\oplus\mathbb{R}^{4k},\,\mathfrak{g}\oplus\mathbb{R}^{4k})]=\mathfrak{g}^{1}\oplus\rho(\mathfrak{g})\,\mathbb{R}^{4k}.$
To prove the second statement, we recall that $\mathfrak{h}$ is $r$-step nilpotent if and only if $r$ is the smallest integer such that, for any $X_{i}\in\mathfrak{g}$ and $V_{i}\in\mathbb{R}^{4k}$,
$\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}=0.$
We first observe that, for any $s\geq 1$,
$\left(\prod_{i=1}^{s}\operatorname{ad}_{(X_{i},0)}\right)(Y,W)=\left(\left(\prod_{i=1}^{s}\operatorname{ad}_{X_{i}}\right)Y,\;\left(\prod_{i=1}^{s}\rho(X_{i})\right)W\right).$
Since $\mathfrak{g}$ is $l$-step nilpotent and $m_{\rho}$ is minimal with the property above, this immediately implies that
$r\geq\mathrm{max}\{l,m_{\rho}\}.$
On the other hand, let $r=\max\{l,m_{\rho}\}$. Let $Y\in\mathfrak{g}$ and $W\in\mathbb{R}^{4k}$. Then, using the fact that
$\left(\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}\right)(Y,W)\in\mathfrak{g}^{r}\oplus\rho(\mathfrak{g})^{\,r}(\mathbb{R}^{4k}),$
and the definition of $l$ and $m_{\rho}$, we conclude that
$\prod_{i=1}^{r}\operatorname{ad}_{(X_{i},V_{i})}=0.$
Therefore, $\mathfrak{h}$ is $r$-step nilpotent with $r=\max\{l,m_{\rho}\}$. ∎ New examples of Obata non flat 2-step nilpotent hypercomplex Lie algebras can be obtained by applying the semidirect product construction, as we show next.
@@
We note that the simply connected Lie groups corresponding to the Lie algebras from examples 4.6 and 4.7 admit lattices, since all the bases considered in the examples are rational.
@@ 5 $\mathbb{H}$-solvability of $2$-step nilpotent Lie algebras
Let $\mathfrak{g}$ be a Lie algebra endowed with a hypercomplex structure $\{J_{1},J_{2},J_{3}\}$. Given a vector subspace $W$ of $\mathfrak{g}$, we denote by $\mathbb{H}W$ the $\{J_{\alpha}\}$-invariant subspace $W+J_{1}W+J_{2}W+J_{3}W$.
Inspired by [Gor1, Definition 1.5], we introduce the following definition.
Definition 5.1.
Let $\mathfrak{g}$ be a Lie algebra endowed with a hypercomplex structure $\{J_{1},J_{2},J_{3}\}$. Define inductively $\{J_{\alpha}\}$-invariant subspaces:
$\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g},\qquad\mathfrak{g}_{k}^{\mathbb{H}}:=\mathbb{H}\mathfrak{g}_{k-1}^{\mathbb{H}},\quad k\geq 1.$
A hypercomplex Lie algebra $(\mathfrak{g},\{J_{\alpha}\})$ is called $\mathbb{H}$-solvable if the following sequence eventually becomes zero for some $k\in\mathbb{N}$:
@@
Let $(\mathfrak{g},\{J_{\alpha}\})$ be an arbitrary hypercomplex Lie algebra.
  1. $(1)$
    For any $k\in\mathbb{N}$, $\mathfrak{g}_{k}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$ and it is an ideal in $\mathfrak{g}_{k-1}^{\mathbb{H}}$.
  2. $(2)$
    If $(\mathfrak{g},\{J_{\alpha}\})$ is $\mathbb{H}$-solvable then $\mathfrak{g}$ is solvable.
We prove the first item. It suffices to show that $\mathfrak{g}_{k}^{\mathbb{H}}$ is an ideal of $\mathfrak{g}_{k-1}^{\mathbb{H}}$. Since $\mathfrak{g}^{1}\subset\mathfrak{g}_{1}^{\mathbb{H}}$, we have that $\mathfrak{g}_{1}^{\mathbb{H}}$ is clearly an ideal of $\mathfrak{g}_{0}^{\mathbb{H}}=\mathfrak{g}$. Assume by induction that $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is an ideal of $\mathfrak{g}_{k-2}^{\mathbb{H}}$. In particular, $\mathfrak{g}_{k-1}^{\mathbb{H}}$ is a subalgebra of $\mathfrak{g}$. Then
@@ 2. $(2)$
To show the second, we work by induction by showing that $\mathfrak{g}_{k}^{\mathbb{H}}\supset D^{k}$, where $\{D^{k}\}_{k\geq 1}$ is the derived series of $\mathfrak{g}$.
For $k=1$, $\mathfrak{g}_{1}^{\mathbb{H}}\supset\mathfrak{g}^{1}=D^{1}$. Assume that $\mathfrak{g}_{k-1}^{\mathbb{H}}\supset D^{k-1}$. Then
$\mathfrak{g}_{k}^{\mathbb{H}}=\mathbb{H}[\mathfrak{g}_{k-1}^{\mathbb{H}},\mathfrak{g}_{k-1}^{\mathbb{H}}]\supset\mathbb{H}[D^{k-1},D^{k-1}]=\mathbb{H}D^{k}\supset D^{k}.$
Therefore, the $\mathbb{H}$-solvability of $(\mathfrak{g},\{J_{\alpha}\})$ implies the solvability of $\mathfrak{g}$. ∎ Let $(\mathfrak{g},J)$ be a nilpotent Lie algebra endowed with a complex structure $J$. It was shown by Salamon [Sal, Corollary 1.4] that the ideal $\mathfrak{g}^{1}+J\mathfrak{g}^{1}$ is a proper ideal of $\mathfrak{g}$. This naturally raises the question of whether an analogous property holds for nilpotent Lie algebras endowed with a hypercomplex structure.
@@
The conjecture holds for Lie algebras of hypercomplex nilmanifolds with flat Obata connection [Gor2], as well as for Lie algebras of abelian hypercomplex nilmanifolds [AV]. The filtration (5.1) corresponds to an iterated hypercomplex toric bundle [Gor1], and in the latter case it was shown in [AV, Proposition 4.5] that such iterated torus bundle terminates at a point. Moreover, for a general complex structure $J_{y}$, the associated complex nilmanifold $(\Gamma\backslash G,J_{y})$ not only admits no complex curves (as a consequence of [Gor1]), but in fact all of its subvarieties are trianalytic.
Here we prove the conjecture for $2$-step nilpotent Lie algebras.
@@
Using that $\mathfrak{g}^{1}\subset\mathfrak{z}$, and Lemma 2.5 (3), we obtain that $J_{\alpha}\mathfrak{g}^{1},J_{\alpha}\mathfrak{g}^{1}]=0$, and
@@
Since by Equation (2.5) we have that $J_{\alpha}[J_{\alpha}\mathfrak{g}^{1},J_{\beta}\mathfrak{g}^{1}]\subset\mathfrak{g}^{1}$, we may write:
@@
We claim that $J_{1}[J_{2}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{2}[J_{1}\mathfrak{g}^{1},J_{3}\mathfrak{g}^{1}]=J_{3}[J_{1}\mathfrak{g}^{1},J_{2}\mathfrak{g}^{1}]$. Using the quaternionic identities and Equation (2.5) with $J_{1}$, we have:
@@ 6 Holonomy of 2-step hypercomplex nilmanifolds
We prove a structural result about the holonomy algebra of the Obata connection on $2$-step nilpotent Lie groups endowed with invariant hypercomplex structures.
@@
Moreover, $A\cdot B=0$ for any $A,B\in\mathfrak{hol}(\nabla)$, therefore, $\mathfrak{hol}(\nabla)$ is a commutative associative algebra with trivial product.
@@
      $J_{2}[\cdot,J_{2}[\cdot,\cdot]]=[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, where the equality follows from Equation (2.5),
@@
      $J_{3}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=J_{1}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows from Equation (2.5),
@@
      $J_{1}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}J_{2}[\cdot,J_{2}[\cdot,\cdot]]=-J_{3}[J_{2}\cdot,J_{2}[\cdot,\cdot]]\in J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$, where the first equality follows by the quaternionic identities and the second equality follows again from Equation (2.5).
    We set $\mathfrak{h}:=\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}+J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. It is not difficult to note that $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant and $\mathfrak{h}\subset\mathfrak{g}_{1}^{\mathbb{H}}\subsetneq\mathfrak{g}$.
@@
      $[Y,X]+J_{1}[J_{1}Y,X]$ $=[Y,X_{2}]+J_{1}[J_{1}Y,X_{2}]$
      $=[Y,X_{2}]-[Y,X_{2}]+[J_{1}Y,J_{1}X_{2}]-J_{1}[Y,J_{1}X_{2}]$
      $=0$
      since $J_{1}X_{2}\in\mathfrak{g}^{1}\subset\mathfrak{z}$, where we have used the integrability of $J_{1}$ in the second equality. ∎
@@
    where we used Lemma 6.2 twice and the $\{J_{\alpha}\}$-invariance of $\mathfrak{h}$.
  • Step 3: Expression for the curvature endomorphisms. Using Steps 1 and 2, we conclude that for any $X,Y\in\mathfrak{g}$,
    $R(X,Y)\in\bigl\{A\in\mathfrak{gl}(n,\mathbb{H})\mid\operatorname{Im}(A)\subset\mathfrak{h}\subset\ker(A)\bigr\}.$
    Moreover, since $R(X,Y)^{2}=0$, we have that $R(X,Y)\in\mathfrak{sl}(n,\mathbb{H})$ and therefore
    $R(X,Y)\in\bigl\{A\in\mathfrak{sl}(n,\mathbb{H})\mid\operatorname{Im}(A)\subset\mathfrak{h}\subset\ker(A)\bigr\}=:\mathfrak{a}.$
@@
    We recall the following version of the Ambrose-Singer Theorem for left-invariant connections.
@@
    [Al] Let $\nabla$ be a left-invariant affine connection on the Lie group $G$, and let $\mathfrak{g}$ denote the Lie algebra of $G$. Then the holonomy algebra $\mathfrak{hol}(\nabla)$ based at the identity element $e\in G$ is the smallest subalgebra of $\mathfrak{gl}(\mathfrak{g})$ containing the curvature endomorphisms $R(X,Y)$ for any $X,Y\in\mathfrak{g}$, and closed under commutators with the left multiplication operators $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$.
    Let us consider $\text{span}\{R(X,Y)\mid X,Y\in\mathfrak{g}\}$. As already observed in Step $3$,
    $\text{span}\{R(X,Y)\mid X,Y\in\mathfrak{g}\}\subset\mathfrak{a}.$
    We point out that $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
    Let $A\in\mathfrak{a}$ and let $X\in\mathfrak{g}$.
@@
    In fact, let $Y\in\mathfrak{g}$ and $Y^{\prime}\in\mathfrak{h}$:
    $A\nabla_{X}Y-\nabla_{X}AY\subset A(\mathfrak{g}_{1}^{\mathbb{H}})-\nabla_{X}\mathfrak{h}=A(\mathfrak{g}_{1}^{\mathbb{H}})\subset\mathfrak{h},$
    that is, $\mathrm{Im}([A,\nabla_{X}])\subset\mathfrak{h}$, and
    $A\nabla_{X}Y^{\prime}-\nabla_{X}AY^{\prime}=0,$
    implying that $\mathfrak{h}\subset\ker([A,\nabla_{X}])$. Moreover, $A,\nabla_{X}]\in\mathfrak{sl}(n,\mathbb{H})$ since $\mathfrak{sl}(n,\mathbb{H})$ is an ideal of $\mathfrak{gl}(n,\mathbb{H})$.
    Since $\mathfrak{a}$ is closed under commutators with $\nabla_{\mathfrak{g}}$, by Theorem 6.3 we get that $\mathfrak{hol}(\nabla)\subset\mathfrak{a}$. In particular, $\mathfrak{a}$ is abelian and has trivial product.
@@
Let $(\mathfrak{g},\{J_{\alpha}\})$ be the $2$-step nilpotent hypercomplex Lie algebra constructed in Example 4.3 and let us consider the corresponding $2$-step nilpotent hypercomplex Lie group $G$. As already observed, the Obata connection $\nabla$ on $G$ is not flat. A direct computation shows that the holonomy algebra $\mathfrak{hol}(\nabla)$ is given by